# BAGELS: Benchmarking the Automated Generation and Extraction of Limitations from Scholarly Text

### v1

@@ A.2 Performance Measurement
Ground Truth OpenReview: Reviewer 1: This paper tries to tackle the domain generalization problem by finding flat minima. It is interesting to investigate the loss landscape of the DNNs on domain generalization problems. Originality: The method proposed in this paper is mainly built on previous approaches such as SWA, and the authors mention this in the submission. Quality: This submission is complete piece of work and the claims are well supported by the empirical results. The evaluation is reasonable. Clarity: This submission is well-written and easy to follow. Significance: This paper provides interesting empirical observations and improves the state-of-the-art by applying their proposed new approach. Pros: The proposed method consistently improves the model performance on domain generalization problems, and the proposed method is flexible and can be easily integrated with other training approaches for better performance. Cons: The generalization bound does not provide intuition for robust risk minimization.. Reviewer 2: Originality The paper is, perhaps by design, not particularly original. Rather, it makes the observation that an existing set of techniques with particular goal also has applications to different goal. This is still, to the best of my knowledge, novel observation and contribution, and furthermore the modifications made to SWA to arrive at SWAD also appear novel. Nevertheless, these modifications are relatively small, so the method can still be viewed as version of the general idea of SWA, and thus originality is not the paper’ strong suit. Quality The paper is of relatively high quality, primarily due to the empirical results. did not carefully check the theory for correctness, but am unsure as to the significance of the theorems. They appear rather similar to prior theoretical results on how flat minima can lead to better generalization, but with some additional divergence terms thrown in as we are dealing with the DG setting… Reviewer 3: Pros: The paper tackles an important and relevant problem to the ML community. The proposed algorithm is relatively simple to implement in practice and perform well experimentally. Cons: My major concern is wrt to the novelty of the work and the connection/motivation between the proposed algorithm SWAD, domain generalization and the derived theoretical insights… '''   LLM Refined OpenReview: "1. . **Connection Between Theory and Method**: The relationship between the theoretical framework presented in the paper and the SWAD algorithm is weak. There is a need for a clearer connection between the flatness term in the equations and the algorithm’s performance.", ’2. . **Novelty of Theoretical Results**: The novelty of the theorem compared to classic results, such as those from Ben-David, is questioned. The theoretical contributions do not significantly advance the understanding of flat minima in the context of domain generalization.’, ’3. . **Generalization Bound Intuition**: The generalization bound provided does not offer sufficient intuition for robust risk minimization (RRM), and it does not convincingly demonstrate that RRM is superior to standard empirical risk minimization (ERM).’], ’4. . **Dependence on Previous Work**: The proposed method heavily relies on the existing SWA approach, with only minor modifications. This raises concerns about the originality of the contribution.’, ’5. . **Theoretical Framework Limitations**: The reliance on VC dimension in the theoretical framework is problematic, as it can be excessively large for the models used, especially in domain generalization scenarios.’, ’6. . **Overlooked Details in Proofs**: There are concerns that important terms have been overlooked in the technical proofs, which could undermine the validity of the results.’, ’7. . **Lack of Clear Motivation for SWAD**: The motivation for using SWAD over SWA is not clearly articulated, making it difficult to understand the advantages of the proposed algorithm.’, ’8. . **Introduction of New Hyperparameters**: The proposed algorithm introduces additional hyperparameters, which may complicate the implementation and tuning process.’, ’9. . **Limited Societal Impact Discussion**: While a limitations section is included, there is a lack of discussion regarding the societal impacts of the research, although specific concerns were not identified.’, "10. . **Potential for Broader Applications**: The method’s applicability beyond domain generalization is not explored, which could limit its significance and impact in the broader machine learning community.
Figure 7: Ground Truth and LLM Extracted Limitations in OpenReview.

### v2

@@ A.2 Performance Measurement
OpenReview: Reviewer 1: This paper tries to tackle the domain generalization problem by finding flat minima. It is interesting to investigate the loss landscape of the DNNs on domain generalization problems. Originality: The method proposed in this paper is mainly built on previous approaches such as SWA, and the authors mention this in the submission. Quality: This submission is complete piece of work and the claims are well supported by the empirical results. The evaluation is reasonable. Clarity: This submission is well-written and easy to follow. Significance: This paper provides interesting empirical observations and improves the state-of-the-art by applying their proposed new approach. Pros: The proposed method consistently improves the model performance on domain generalization problems, and the proposed method is flexible and can be easily integrated with other training approaches for better performance. Cons: The generalization bound does not provide intuition for robust risk minimization.. Reviewer 2: Originality The paper is, perhaps by design, not particularly original. Rather, it makes the observation that an existing set of techniques with particular goal also has applications to different goal. This is still, to the best of my knowledge, novel observation and contribution, and furthermore the modifications made to SWA to arrive at SWAD also appear novel. Nevertheless, these modifications are relatively small, so the method can still be viewed as version of the general idea of SWA, and thus originality is not the paper’ strong suit. Quality The paper is of relatively high quality, primarily due to the empirical results. did not carefully check the theory for correctness, but am unsure as to the significance of the theorems. They appear rather similar to prior theoretical results on how flat minima can lead to better generalization, but with some additional divergence terms thrown in as we are dealing with the DG setting… Reviewer 3: Pros: The paper tackles an important and relevant problem to the ML community. The proposed algorithm is relatively simple to implement in practice and perform well experimentally. Cons: My major concern is wrt to the novelty of the work and the connection/motivation between the proposed algorithm SWAD, domain generalization and the derived theoretical insights… '''  LLM extracted OpenReview: "1. . **Connection Between Theory and Method**: The relationship between the theoretical framework presented in the paper and the SWAD algorithm is weak. There is a need for a clearer connection between the flatness term in the equations and the algorithm’s performance.", ’2. . **Novelty of Theoretical Results**: The novelty of the theorem compared to classic results, such as those from Ben-David, is questioned. The theoretical contributions do not significantly advance the understanding of flat minima in the context of domain generalization.’, ’3. . **Generalization Bound Intuition**: The generalization bound provided does not offer sufficient intuition for robust risk minimization (RRM), and it does not convincingly demonstrate that RRM is superior to standard empirical risk minimization (ERM).’], ’4. . **Dependence on Previous Work**: The proposed method heavily relies on the existing SWA approach, with only minor modifications. This raises concerns about the originality of the contribution.’, ’5. . **Theoretical Framework Limitations**: The reliance on VC dimension in the theoretical framework is problematic, as it can be excessively large for the models used, especially in domain generalization scenarios.’, ’6. . **Overlooked Details in Proofs**: There are concerns that important terms have been overlooked in the technical proofs, which could undermine the validity of the results.’, ’7. . **Lack of Clear Motivation for SWAD**: The motivation for using SWAD over SWA is not clearly articulated, making it difficult to understand the advantages of the proposed algorithm.’, ’8. . **Introduction of New Hyperparameters**: The proposed algorithm introduces additional hyperparameters, which may complicate the implementation and tuning process.’, ’9. . **Limited Societal Impact Discussion**: While a limitations section is included, there is a lack of discussion regarding the societal impacts of the research, although specific concerns were not identified.’, "10. . **Potential for Broader Applications**: The method’s applicability beyond domain generalization is not explored, which could limit its significance and impact in the broader machine learning community.
Figure 8: Tool extracted OpenReview and LLM Extracted OpenReview.